# Semantic Mapping in Indoor Embodied AI – A Comprehensive Survey and Future Directions

### v1

@@ 4.3 Point-cloud map
There are typically two strategies to train neural fields 1) use distillation (Peng et al., 2023; Kerr et al., 2023; Taioli et al., 2023; Qin et al., 2024; Guo et al., 2024; Qiu et al., 2024) to provide features that are similar to a pretrained 2D backbone, such as CLIP (Radford et al., 2021), 2) use of differentiable renderer to match the rendered semantics in addition to color (Zhi et al., 2021; Vora et al., 2021).
OpenScene (Peng et al., 2023), for example, predicts dense 3D features so that they are co-embedded with the corresponding text and the image in the CLIP embedding space. This allows the association of each 3D point in the scene with semantic information such that the scene can be queried using text to infer physical properties, affordances, etc. CLIP2Scene (Chen et al., 2023c) also uses CLIP to perform a 3D point cloud segmentation on outdoor scenes for application in autonomous driving. However, because the optimization happens per scene in many of these methods, they are typically expensive and not suitable for real-time use in embodied applications. A representative work that uses neural field maps for embodied AI include CLIP-Fields (Shafiullah et al., 2023). Zhang et al. (2023a), on the other hand, builds a 3D semantic scene representation based on an online point cloud-based construction algorithm (Zhang et al., 2020) made efficient by using a tree-based dynamic data structure. This method is very memory efficient and a lot faster when applied to the ObjectNav task. Similarly, Lei et al. (2024) shows that a 3D representation based on Gaussian Splatting is able to achieve state-of-the-art performance on the ImageNav task.
Summary. Use of such dense map remain relatively unexplored in the context of embodied AI tasks and can be a promising a future direction. However, using such a dense map could lead to significant memory consumption and computational inefficiency during querying. Moreover, in a typical indoor scene, most of the 3D space is empty and might not be useful in reasoning about the environment.
4.4 Hybrid map
@@
So far we have seen how prior works structure maps as either a spatial grid or a landmark based scene graph or a more dense point-cloud. However there is a more recent effort on combining these different structures into a single map representation. This helps to capture information at various granularity and perform different types of reasoning on the environment.
@@
The combination of metric information from grid-based maps together with a topological map is also known as a topometric map (Thrun et al., 1998; Tomatis et al., 2001; Blanco et al., 2008; Konolige et al., 2011; Ko et al., 2013; An et al., 2023). Thrun et al. (1998) proposes a single statistical mapping algorithm that first constructs a coarse topological map and uses it to construct a fine-grained grid map. They show that the topological map solves a global alignment problem by correcting large odometry errors, while the grid map solves a local alignment problem by producing high-resolution maps. Tomatis et al. (2001), on the other hand, proposes a compact environmental model where corners and hallways are represented by a topological map and rooms are represented by a grid map, both of which are connected in a single representation. When the robot is moving in hallways, it creates and updates the global topological map, and as soon as it enters a room, it creates a new local metric map [^1] They differentiate hallways and rooms by using laser sensor such that thin long open spaces are considered as hallways whereas other open spaces are considered as rooms.. They argue that the robot will only need to be precise inside rooms (e.g. manipulating objects, etc.) which justifies the need for the fine-grained precise metric maps, while the topological map is used to simply maintain global consistency in indistinguishable spaces such as long hallways and transitioning between significant places. BEVBert (An et al., 2023) is a more recent method that constructs hybrid maps offline and then learn a multimodal map representation to perform better spatial reasoning in the complex task of language-guided navigation.
@@
There are also hybrid maps that combines grids, point-clouds, and topological maps. StructNav (Chen et al., 2023b) builds such a hybrid map where the spatial grid stores occupancy information, a scene graph stores landmarks with their connectivity, and a 3D semantic point cloud where each 3D point in the environment has a semantic label associated with it. Thus the spatial grid allows for obstacle avoidance and low-level path planning, the scene graph allows high-level reasoning about the relationship among the landmarks, and the 3D point cloud allows for a more dense semantic and spatial matching.
@@
It is also possible to build maps that capture the semantic hierarchy of scenes at various levels of abstraction that allows for various levels of reasoning. Armeni et al. (2019) represent a static scene at multiple levels of hierarchy (buildings, rooms and objects), where entities are represented as nodes in a hierarchical graph and connected by edges representing coordinate frame transformations. Rosinol et al. (2020a) (Dynamic Scene Graphs) extend this to dynamic scenes by having layered scene graphs in addition to a dense 3D metric spatial map and modeling spatio-temporal relations between objects and agents. Hughes et al. (2024) demonstrate that hierarchical scene representations scale better than flat representations in a large environments and thereafter introduces a system called Hydra that incrementally builds 3D scene graphs from sensor data in real-time. Fischer et al. (2024) introduces a multi-level scene graph representation of large-scale dynamic urban environments from a set of images captured from moving vehicles and proposes a new view synthesis benchmark for urban driving scenarios.
@@
Summary. Different types of map structures have their own strengths and limitations. While the coarse topological map is useful to represent significant landmarks in an environment, fine-grained metric maps are useful to represent its precise geometry and point-cloud map is useful to represent an even more dense 3D geometry of objects in the scene. Therefore it’s crucial to combine two or more of these structures in order to create better representations of the environment, preferably at varying levels of semantic abstraction. While hybrid maps have been explored in robotics, they still remain under-explored in embodied AI. However, weaknesses of these maps need to be considered carefully before combining them. For example, topological and point-cloud map scale better than a grid map with larger environments, and point-cloud map needs the most and topological map needs the least memory to be stored.
@@ 5.1 Explicit encoding
Many prior works store explicit information in the map. Prior works have found that almost every embodied AI task benefits from the information about obstacles present in the environment and as such maintaining a spatial map that has this occupancy information helps the agent learn to avoid obstacles on its path. In such cases, the spatial map stores a binary occupancy value of $1$ or 00 depending on whether the corresponding location in the physical environment is ‘occupied’ by objects or not.
@@
In an exploration task, however, the agent needs to maximize the explored area in an environment while being efficient, in which cases, the information of whether a location has already been explored encourages the agent to explore unexplored areas more. Active Neural SLAM  (Chaplot et al., 2019) stores the explored (binary) information in addition to the occupancy information. Active Neural SLAM consists of several modules connected together to perform the task of Exploration to maximize the explored area. The ‘Neural SLAM’ module takes as inputs the visual observations and agent pose and outputs a top-down egocentric spatial map by learning to predict occupancy and explored information using a binary cross-entropy loss. All the modules are jointly trained with different losses.
@@
However, in more complex tasks such as ObjectNav, which requires the agent to navigate to a particular semantic category in its environment, storing occupancy and explored information alone might not be sufficient. In such cases, it is important for the agent to identify the semantic category of the object. Goal-oriented Semantic Exploration (SemExp) by Chaplot et al. (2020a) additionally stores the semantic class labels of the objects that the agent identifies through its visual observations. SemExp uses a MaskRCNN (He et al., 2017) on the RGB observations to predict the semantic categories of the objects and then project these on the map using depth observations. It aggregates the occupancy and explored information using element-wise max pooling whereas the semantic categories are overwritten by the latest prediction. A denoising network is then used to get the final map. This work demonstrates that using this semantic map to predict a long-term goal helps the agent find the goal object category more efficiently. The mapping module is trained using supervised learning with cross-entropy loss on the predicted semantic map. Similarly, a semantic categories map also helps a more complex longer-horizon task of MultiON. MOPA (Raychaudhuri et al., 2023) shows that maintaining a memory with objects that the agent observes while moving around is crucial to perform this task efficiently. However, a map built by segmenting the image first and then projecting may result in ‘label splattering’ i.e. noisy category labels splattered across multiple grid cells in a spatial map. This arises mostly due to noisy depth observations and might negatively affect agent performance in a task. Semantic MapNet (Cartillier et al., 2021) finds that projecting encoded features on to a map and then segmenting produces a more noise-free map. They show that this map can be then be applied to two different tasks effectively. However this method requires an additional exploration phase where the agent effectively explores the environment to build the map first and then use that map in the downstream task. A recent work GOAT (Chang et al., 2023; Khanna et al., 2024) shows that having an object instance map helps to navigate to a goal specified by either language, image or a category label. They achieve this by storing raw images and later using CLIP (Khandelwal et al., 2022) for image-to-language matching and SuperGlue (Sarlin et al., 2020) for image-to-image matching.
@@
While the above works build explicit spatial maps, some works also build explicit topological maps. Kwon et al. (2021) store the visitation timestep in the graph node in order to encode a temporal relation between visited locations. This information is then replaced by the latest visitation timestep while aggregating the map.
@@ 5.2.1 Closed-vocabulary encoding
ConceptGraphs (2023) AI2Thor, robot CLIP, DINO GPT-4 highest similarity score
@@ 5.2.2 Open-vocabulary encoding
The limitation of the closed-vocabulary encoding can be mitigated by extracting features from a Large Vision-Language Model (LVLM) that was jointly trained on a vast amount of internet data of images and their text captions, such as CLIP (Radford et al., 2021). This allows the map to store information about ‘any’ object in the environment and eventually be queried via an open-vocabulary text query not limited to a predefined set of object categories. For example, an agent may be asked to ‘find a red and blue striped zebra toy in the children’s room’. It is likely that the agent has not seen a ‘red and blue striped zebra toy’ during training but can leverage a LVLM to reason about its prior knowledge about zebras, colors and rooms in general.
Moreover, recently large language models (LLMs), such as GPT-4 (OpenAI, 2023), have been shown to be able to perform complex task planning. Thus open-vocabulary maps built using LVLM along with LLM-based planners have led to a recent line of works in embodied agents (Tab. 5, Fig. 8).
A pretrained CLIP model can be used to compute similarity scores between an input image and a natural language description with the highest score corresponding to the most likely image-text match. CLIP has been successfully used in map building where the map stores the similarity scores between each image that the agent observes and the language instruction describing an object. Popularly, these maps are called value maps. CoW (Gadre et al., 2023) shows that such a 2D value map can be successfully applied in the downstream task of language-driven ObjectNav in a zero-shot manner without any re-training. The planner in this method plans a path to the object when the stored similarity score exceeds a certain threshold. VLFM (Yokoyama et al., 2023) follows a similar strategy to perform ObjectNav task by using the BLIP-2 (Li et al., 2023) 2D value map to semantically explore the environment. VoxPoser (Huang et al., 2023b), on the other hand, builds a similar 3D value map to efficiently perform table-top robot manipulation. Such map encoding is quite powerful compared to the previous encodings that stores the semantic labels for a predefined set of objects. However, this particular map still lacks the ability to perform spatial reasoning because it performs similarity matching to the entire input image ignoring the semantic information about individual objects in the scene.
Figure 8: Open-Vocabulary map building. There has been a growing interest to build flexible open-vocabulary maps which can be built once and then used in various downstream tasks during inference. VLMap (Huang et al., 2023a) and NLMap (Chen et al., 2023a) structure their maps as spatial maps while ConceptGraphs (Gu et al., 2023) build a topological map. (Figures reproduced from paper.)
@@
While NLMap builds a 3D spatial map, ConceptGraphs (Gu et al., 2023) build a topological map following a similar approach. It retrieves the objects of interest from the input image by using the class-agnostic 2D segmentation model, Segment-Anything (SAM) (Kirillov et al., 2023) to obtain candidate masks. These objects then form the nodes of the topological map. The image features for each object can then be extracted by CLIP and DINO (Oquab et al., 2023) to be stored in the corresponding node. Additional information about the objects can also be stored in each node. For example, ConceptGraphs store the point cloud of the proposed masks and a caption for each object as obtained by using LLaVA (Liu et al., 2023a) and GPT-4 (OpenAI, 2023) along with a point cloud obtained by projecting the object mask proposed by SAM into 3D space. After the nodes are formed, edges between the nodes can be constructed depending on whether the objects are spatially related. In ConceptGraphs, spatial relation between two objects is determined by whether the point clouds of the respective objects have a geometric similarity or overlap. In other words, when a certain proportion of points in point cloud of one object lie within a distance threshold of that of the second object, the objects can be said to be spatially related to each other and an edge is constructed between the corresponding nodes. The edges additionally store a spatial relationship description obtained from LLM. When a newly detected object is found to be similar to a node, it is updated with averaged object features, union of point clouds and the latest object caption. In case it is not similar, a new node is added to the graph. ConceptGraphs show that a single map representation built in this manner can be successfully and effectively used in object grounding, robot navigation and robot manipulation tasks.

### v2

@@ 4.3.2 Neural fields
Neural fields are continuous functions that map spatial coordinates (and sometimes viewing directions) to signals such as color, occupancy or semantic features. Unlike point clouds that discretely store information, neural fields encode entire 3D scenes as compact, continuous functions. Often, a neural network (typically MLP) is used to obtain the features associated with each point given the position (x,y,z) as input (Mildenhall et al., 2021; Mescheder et al., 2019). There are typically two strategies to train neural fields 1) use distillation (Kerr et al., 2023; Qiu et al., 2024) to provide features that are similar to a pretrained 2D backbone, such as CLIP (Radford et al., 2021), and 2) use of differentiable renderer to match the rendered semantics in addition to color (Zhi et al., 2021; Vora et al., 2021).
In robotics, recent works (Sucar et al., 2021; Zhu et al., 2022; Rosinol et al., 2023) have integrated neural fields into SLAM systems for real-time mapping. While these enable compact, high-fidelity 3D reconstruction and view synthesis, they are more computationally intensive and less interpretable than traditional methods. Thus challenges remain in training speed, memory demands and dynamic scene handling. In contrast, neural Signed Distance Fields (SDFs) are a specific type of neural field that encode the signed distance to the closest surface, enabling precise geometric reasoning. Neural SDFs have been shown to be particularly suited for tasks like mapping, collision checking, and planning due to their ability to represent detailed geometry with differentiable distance metrics (Ortiz et al., 2022; Camps et al., 2022; Talha Bukhari et al., 2025). In embodied AI, a representative work using neural field maps include CLIP-Fields (Shafiullah et al., 2023), which learns a continuous, neural implicit representation of 3D environments by aligning 2D image features from CLIP with 3D points in space using volumetric rendering. This enables the agent to build an open-vocabulary, queryable 3D map where each spatial point can be semantically queried using natural language, without needing task-specific supervision.
In 3D scene understanding, neural fields have emerged as a powerful representation from early works on NeRF (Mildenhall et al., 2021) to incorporating semantics into neural fields, enabling tasks such as semantic segmentation, object discovery and scene parsing in 3D (Fu et al., 2022). More recent advances (Wang et al., 2022; Kerr et al., 2023) introduce language-supervised neural fields by integrating CLIP features, enabling open-vocabulary semantic reasoning and text-driven queries over 3D scenes.
Summary. Neural fields are gaining popularity in robotics and embodied AI for building continuous 3D scene representations that capture both geometry and semantics. However, challenges remain in real-time training and deployment in dynamic environments. An important future direction would be to focus on making neural fields more efficient, adaptable, and compatible with foundation models for open-world semantic reasoning.
4.4 Hybrid maps
@@
So far we have seen how prior works structure maps as either a spatial grid or a landmark based scene graph or a more dense geometric map. However there is a more recent effort on combining these different structures into a single map representation. This helps to capture information at various granularity and perform different types of reasoning on the environment.
@@
The combination of metric information from grid-based maps together with a topological map is also known as a topometric map (Thrun et al., 1998b; Tomatis et al., 2001; Blanco et al., 2008b; Konolige et al., 2011; Ko et al., 2013; An et al., 2023) (Mozos et al., 2007; Vasudevan & Siegwart, 2008; Zender et al., 2008; Pronobis et al., 2010; Hemachandra et al., 2011; Walter et al., 2013; 2014; Hemachandra et al., 2015; Duvallet et al., 2016; Walter et al., 2022). Thrun et al. (1998b) proposes a single statistical mapping algorithm that first constructs a coarse topological map and uses it to construct a fine-grained grid map. They show that the topological map solves a global alignment problem by correcting large odometry errors, while the grid map solves a local alignment problem by producing high-resolution maps. Tomatis et al. (2001), on the other hand, proposes a compact environmental model where corners and hallways are represented by a topological map and rooms are represented by a grid map, both of which are connected in a single representation. When the robot is moving in hallways, it creates and updates the global topological map, and as soon as it enters a room, it creates a new local metric map [^1] They differentiate hallways and rooms by using laser sensor such that thin long open spaces are considered as hallways whereas other open spaces are considered as rooms.. They argue that the robot will only need to be precise inside rooms (e.g. manipulating objects, etc.) which justifies the need for the fine-grained precise metric maps, while the topological map is used to simply maintain global consistency in indistinguishable spaces such as long hallways and transitioning between significant places. BEVBert (An et al., 2023) is a more recent method that constructs hybrid maps offline and then learn a multimodal map representation to perform better spatial reasoning in the complex task of language-guided navigation.
@@
There are also hybrid maps that combines grids, dense geometric, and topological maps. StructNav (Chen et al., 2023b) builds such a hybrid map where the spatial grid stores occupancy information, a scene graph stores landmarks with their connectivity, and a 3D semantic point cloud where each 3D point in the environment has a semantic label associated with it. Thus the spatial grid allows for obstacle avoidance and low-level path planning, the scene graph allows high-level reasoning about the relationship among the landmarks, and the 3D point cloud allows for a more dense semantic and spatial matching.
@@
It is also possible to build maps that capture the semantic hierarchy of scenes at various levels of abstraction (Galindo et al., 2005; Pronobis & Jensfelt, 2012; Hemachandra et al., 2014; Werby et al., 2024) that allows for various levels of reasoning. Armeni et al. (2019) represent a static scene at multiple levels of hierarchy (buildings, rooms and objects), where entities are represented as nodes in a hierarchical graph and connected by edges representing coordinate frame transformations. Tang et al. (2025) (OpenIn) build a similar hierarchical scene graph to track objects in dynamically changing indoor environments. Rosinol et al. (2020a) (Dynamic Scene Graphs) too build layered scene graphs to track moving agents and objects in addition to building a dense 3D metric spatial map by modeling spatio-temporal relations between objects and agents. Hughes et al. (2024) demonstrate that hierarchical scene representations scale better than flat representations in a large environments and thereafter introduces a system called Hydra that incrementally builds 3D scene graphs from sensor data in real-time. Fischer et al. (2024) introduces a multi-level scene graph representation of large-scale dynamic urban environments from a set of images captured from moving vehicles and proposes a new view synthesis benchmark for urban driving scenarios.
@@
Summary. Different types of map structures have their own strengths and limitations. While the coarse topological map is useful to represent significant landmarks in an environment, fine-grained metric maps are useful to represent its precise geometry and dense geometric map is useful to represent an even more dense 3D geometry of objects in the scene. Therefore it’s crucial to combine two or more of these structures in order to create better representations of the environment, preferably at varying levels of semantic abstraction. While hybrid maps have been explored in robotics, they still remain under-explored in embodied AI. However, weaknesses of these maps need to be considered carefully before combining them. For example, topological and dense geometric maps scale better than a grid map with larger environments, and dense geometric map needs the most and topological map needs the least memory to be stored.
@@ 5.1 Explicit encoding
Many prior works store explicit information in the map, such as occupancy, which has proven beneficial in obstacle avoidance (Elfes, 1989; May et al., 2009; Chaplot et al., 2019; Georgakis et al., 2022a). In such cases, the spatial map stores a binary occupancy value of $1$ or 00 depending on whether the corresponding location in the physical environment is ‘occupied’ by objects or not.
@@
In an exploration task, however, the agent needs to maximize the explored area in an environment while being efficient, in which cases, the information of whether a location has already been explored encourages the agent to explore unexplored areas more. Active Neural SLAM (Chaplot et al., 2019) stores the explored (binary) information in addition to the occupancy information. Active Neural SLAM consists of several modules connected together to perform the task of Exploration to maximize the explored area. The ‘Neural SLAM’ module takes as inputs the visual observations and agent pose and outputs a top-down egocentric spatial map by learning to predict occupancy and explored information using a binary cross-entropy loss. All the modules are jointly trained with different losses. VLMNav (Goetting et al., 2024) similarly stores explored information in a top-down voxel map to demonstrate its generalizability to downstream navigation tasks.
@@
However, in more complex tasks such as ObjectNav, which requires the agent to navigate to a particular semantic category in its environment, storing occupancy and explored information alone might not be sufficient. In such cases, it is important for the agent to identify the semantic category of the object (Meger et al., 2008; Georgakis et al., 2021; Chaplot et al., 2021; Raychaudhuri et al., 2023; Khanna et al., 2024; Zhang et al., 2025). Goal-oriented Semantic Exploration (SemExp) by Chaplot et al. (2020a) additionally stores the semantic class labels of the objects that the agent identifies through its visual observations. SemExp uses a MaskRCNN (He et al., 2017a) on the RGB observations to predict the semantic categories of the objects and then project these on the map using depth observations. It aggregates the occupancy and explored information using element-wise max pooling whereas the semantic categories are overwritten by the latest prediction. A denoising network is then used to get the final map. This work demonstrates that using this semantic map to predict a long-term goal helps the agent find the goal object category more efficiently. The mapping module is trained using supervised learning with cross-entropy loss on the predicted semantic map. Similarly, a semantic categories map also helps a more complex longer-horizon task of MultiON. MOPA (Raychaudhuri et al., 2023) shows that maintaining a memory with objects that the agent observes while moving around is crucial to perform this task efficiently. However, a map built by segmenting the image first and then projecting may result in ‘label splattering’ i.e. noisy category labels splattered across multiple grid cells in a spatial map. This arises mostly due to noisy depth observations and might negatively affect agent performance in a task. Semantic MapNet (Cartillier et al., 2021) finds that projecting encoded features on to a map and then segmenting produces a more noise-free map. They show that this map can be then be applied to two different tasks effectively. However this method requires an additional exploration phase where the agent effectively explores the environment to build the map first and then use that map in the downstream task. A recent work GOAT (Chang et al., 2023; Khanna et al., 2024) shows that having an object instance map helps to navigate to a goal specified by either language, image or a category label. They achieve this by storing raw images and later using CLIP (Khandelwal et al., 2022) for image-to-language matching and SuperGlue (Sarlin et al., 2020) for image-to-image matching. MapNav (Zhang et al., 2025) builds an annotated semantic map by storing text labels of the segmented objects and shows that this helps a VLM to ground objects better.
@@
While the above works build explicit spatial maps, some works also build explicit topological maps (Choset & Nagatani, 2001; Blöchliger et al., 2018; Chen et al., 2022; Salas-Moreno et al., 2013; Duvallet et al., 2013; Patki et al., 2019; Arkin et al., 2020a; Raychaudhuri et al., 2025; Chang et al., 2025). Kwon et al. (2021) store the visitation timestep in the graph node in order to encode a temporal relation between visited locations. This information is then replaced by the latest visitation timestep while aggregating the map.
@@ 5.2.1 Closed-vocabulary encoding
ConceptGraphs (2023b) AI2Thor, robot CLIP, DINO GPT-4 highest similarity score
@@
LM-Nav (2023) robot CLIP modified Dijkstra CLIP similarity
ConceptFusion (2023) AI2Thor, robot CLIP LLM-based weighted mean
Le-RNR-Map (2023) Habitat CLIP Random walk exploration latest
HOV-SG (2024) Habitat, robot CLIP A* weighted mean
RoboHop (2024) Habitat, robot CLIP Dijkstra mean
Clio (2024) robot CLIP Dijkstra mean
OVL-Map (2025) Habitat LSeg DD-PPO (2019b) weighted mean
@@ 5.2.2 Open-vocabulary encoding
The limitation of the closed-vocabulary encoding can be mitigated by extracting features from a Large Vision-Language Model (LVLM) that was jointly trained on a vast amount of internet data of images and their text captions, such as CLIP (Radford et al., 2021). This allows the map to store information about ‘any’ object in the environment and eventually be queried via an open-vocabulary text query (Wen et al., 2025) not limited to a predefined set of object categories. For example, an agent may be asked to ‘find a red and blue striped zebra toy in the children’s room’. It is likely that the agent has not seen a ‘red and blue striped zebra toy’ during training but can leverage a LVLM to reason about its prior knowledge about zebras, colors and rooms in general. Moreover, recently large language models (LLMs), such as GPT-4 (OpenAI, 2023), have been shown to be able to perform complex task planning. Thus open-vocabulary maps built using LVLM along with LLM-based planners (Taioli et al., 2023; Huang et al., 2023b; Long et al., 2024) have led to a recent line of works in embodied agents (Tab. 6, Fig. 8).
A pretrained CLIP model can be used to compute similarity scores between an input image and a natural language description with the highest score corresponding to the most likely image-text match. CLIP has been successfully used in map building where the map stores the similarity scores between each image that the agent observes and the language instruction describing an object. Popularly, these maps are called value maps. CoW (Gadre et al., 2023) shows that such a 2D value map can be successfully applied in the downstream task of language-driven ObjectNav in a zero-shot manner without any re-training. The planner in this method plans a path to the object when the stored similarity score exceeds a certain threshold. VLFM (Yokoyama et al., 2023) follows a similar strategy to perform ObjectNav task by using the BLIP-2 (Li et al., 2023) 2D value map to semantically explore the environment. InstructNav (Long et al., 2024) extends this idea to enhance semantic value maps with multi-sourced value maps encoding actions, landmarks, and navigation history, thus improving generic instruction following. VoxPoser (Huang et al., 2023b), on the other hand, builds a similar 3D value map to efficiently perform table-top robot manipulation. Such map encoding is quite powerful compared to the previous encodings that stores the semantic labels for a predefined set of objects. However, this particular map still lacks the ability to perform spatial reasoning because it performs similarity matching to the entire input image ignoring the semantic information about individual objects in the scene.
Figure 8: Open-Vocabulary map building. There has been a growing interest to build flexible open-vocabulary maps which can be built once and then used in various downstream tasks during inference. VLMap (Huang et al., 2023a) and NLMap (Chen et al., 2023a) structure their maps as spatial maps while ConceptGraphs (Gu et al., 2023b) build a topological map. (Figures reproduced from paper.)
@@
While NLMap builds a 3D spatial map, many (Shah et al., 2023; Garg et al., 2024) build a topological map following a similar approach. For instance, ConceptGraphs (Gu et al., 2023b) is a recent approach that retrieves the objects of interest from the input image by using the class-agnostic 2D segmentation model, Segment-Anything (SAM) (Kirillov et al., 2023) to obtain candidate masks. These objects then form the nodes of the topological map. The image features for each object can then be extracted by CLIP and DINO (Oquab et al., 2023) to be stored in the corresponding node. Additional information about the objects can also be stored in each node. For example, ConceptGraphs store the point cloud of the proposed masks and a caption for each object as obtained by using LLaVA (Liu et al., 2023a) and GPT-4 (OpenAI, 2023) along with a point cloud obtained by projecting the object mask proposed by SAM into 3D space. After the nodes are formed, edges between the nodes can be constructed depending on whether the objects are spatially related. In ConceptGraphs, spatial relation between two objects is determined by whether the point clouds of the respective objects have a geometric similarity or overlap. In other words, when a certain proportion of points in point cloud of one object lie within a distance threshold of that of the second object, the objects can be said to be spatially related to each other and an edge is constructed between the corresponding nodes. The edges additionally store a spatial relationship description obtained from LLM. When a newly detected object is found to be similar to a node, it is updated with averaged object features, union of point clouds and the latest object caption. In case it is not similar, a new node is added to the graph. ConceptGraphs show that a single map representation built in this manner can be successfully and effectively used in object grounding, robot navigation and robot manipulation tasks.